We consider the sequence of real numbers (Un) defined on N by Uo = -1, U1 = 1/2 and for every n E N, U(n+2) = U(n+1) - 1/4 Un. Where N reprents the set of natural numbers. Vn = U(n+1) - (1/2)Un.
We define the sequence (Wn) by for every n E N, (Wn) = Un / Vn.
(i) Deduce an expression for Un in terms of n.
(ii) For every natural number n, we set Sn = Uo + U1 + U2 + ...+ Un. Show by induction that for every n E N, Sn = 2 - (2n + 3)/ 2^n.
1
Expert's answer
2021-02-02T04:28:21-0500

Solution


  • To estimate the common term of the sequence, try finding the first few terms with the help of the first relationship given

"\qquad\qquad
\begin{aligned}
n=0 \to u_2&=u_1-\frac{1}{4}u_0\\
u_2&=\frac{1}{2}+\frac{1}{4}=\frac{3}{4}\\
n=1\to u_3&=u_2-\frac{1}{4}u_1\\
u_3&=\frac{3}{4}-\frac{1}{8}=\frac{5}{8}\\

\end{aligned}"

  • Then the sequence can be written as,

"-1,\frac{1}{2},\frac{3}{4},\frac{5}{8},......."

  • It can be rearranged to

"\frac{-1}{1},\frac{1}{2},\frac{3}{4},\frac{5}{8},....."

  • Now treating the numerator & denominator separately, common term for each can be written as follows

"\qquad\qquad
\begin{aligned}
U_r&=\frac{2r-1}{2^r}
\end{aligned}"

  • Then the common term can be written in terms of n as follows by replacing r with n

"\qquad\qquad
\begin{aligned}
U_n&= \frac{2n-1}{2^n}
\end{aligned}"

2)

  • To check with induction, first try the validity for the first possible natural numbers 0 & 1.

"\qquad\qquad
\begin{aligned}
S_n&=2-\frac{(2n+3)}{2^n}\\
n=0\\
S_0&= 2-3=-1=U_0\cdots(\checkmark)\\
n=1\\
S_1&= 2-\frac{5}{2}=\frac{-1}{2}=U_2\cdots(\checkmark)
\end{aligned}"

  • Then assume the equation holds true for an arbitrary integer "p".

"\qquad\qquad
\begin{aligned}
S_p&= 2-\frac{(2p+3)}{2^p}\\
\end{aligned}"

  • Now check whether the equation holds further true for integer "p+1" ,

"\qquad\qquad
\begin{aligned}
S_{p+1}&=S_p+U_{p+1}\\
&=2-\frac{(2p+3)}{2^p}+\frac{2(p+1)-1}{2^{p+1}}\\
&=2-\Bigg[\frac{2(2p+3)-2(p+1)+1}{2^{p+1}}\Bigg]\\
&= 2-\Bigg[\frac{4p+6-2p-2+1}{2^{p+1}}\Bigg]\\
&= 2-\frac{(2p+5)}{2^{p+1}}\\
&\text{rearranging \, in\, terms \,of\,p+1}\\
S_{p+1}&=2-\frac{[2(p+1)+3]}{2^{p+1}}\cdots(\checkmark)
\end{aligned}"

  • As the equation can be further written for the "p+1" term, it can be concluded that the equation holds true for all "n \in N"